Given starting vocabulary

English Spanish
tooth diente
joy alegria
tree arbol
electricity electricidad
table tabla
money dinero
sound sonido
music musica

Unknown-permutation vocabulary

plant bailar
car hablar
dance amigo
speak coche
friend planta